# Print-N-Grip: A Disposable, Compliant, Scalable and One-Shot 3D-Printed Multi-Fingered Robotic Hand

### v1

@@ I Introduction
Figure 2: (a) Front and side view of a PNG finger. Internal view of the proximal link is seen in the circle where the two bands are separated by a bulge. Due to its height, the bulge prevents from the contact pad to press onto the bands. (b) Different scales $\kappa$ of a PNG-2 hand. As convention, a hand with a finger length of 52 mm is defined as $\kappa=1$. (c) Examples of three- (PNG-3) and four- (PNG-4) finger hands. Both hands are one-shot 3D printed. (d) Universal base for mounting the printed hand. The 3D printed hand is fixed to the base by inserting the mounting pins into the conical grooves and closing with the tightening cap. A reduction sleeve is used for hands of small scale. (e) Fingernail on the distal link assists in picking-up small objects.
@@ II Related work
One-shot 3D printing in the field of robotics is not a novel concept. Yet, not much prior work exists on one-shot 3D printing of robotic hands. Of the few, recent work have proposed a one-shot printed granular jamming gripper [21]. Another work focused on one-shot fabrication of a finger while embedding tactile sensors and shape memory alloy actuator [22]. A compliant parallel gripper was proposed and designated for extracting solid objects from powder [23]. Similarly, a bio-inspired parallel gripper was proposed where the joint of each finger receives its compliance from a spiral spring [24]. While innovative, the proposed parallel grippers are not tendon-based and cannot offer high capacity lifting and durability. A more recent work introduced a one-shot 3D printed two-finger gripper which operates on pneumatic pressure [25]. A work most related to ours proposed a one-shot printed hand with an integrated tendon line [26]. Fabrication of the hand requires a high-cost multi jet fusion printer and, due to the thin tendon, cannot provide high-capacity power grasping or scalability. In addition, it does not enable rapid and easy replacement using a universal base. While not one-shot, a recent and intriguing work has proposed the two-finger underactuated InstaGrasp gripper [27]. The gripper has the structure of the OpenHand [12] Model T42 where each finger is composed of nine 3D printed parts of Polylactic Acid (PLA) and TPU. While almost entirely 3D printed and compliant, the hand requires some assembly. Unlike the proposed PNG hand in this paper, these 3D printed robotic hands do not demonstrate high-load grasping, are not scalable and do not offer easy replacement for disposing if contaminated.
@@ III-B Multi-finger hand
A PNG hand is printed with $n$ fingers in one-shot with a standard FDM printer. For simple notation, we denote an $n$-finger PNG hand as PNG-$n$, e.g., PNG-3 denotes a three-finger PNG hand. Thermoplastic Polyurethane (TPU) filament is a flexible and strong material and, thus, used for the 3D printing. Note that TPU has a heat resistance of up to approximately ${}^{\circ}$. Hence, the hand can go through multi-use sterilization by thermal disinfection at ${}^{\circ}$ or Gamma sterilization [33]. We make a distinction between a two-finger (PNG-2) and a three or more finger ($n>2$) hands described next.
@@ III-B1 PNG-2, Two-finger hands
A PNG-2 hand is seen in Figure 2b. The tendon bands of both fingers are connected at a pull loop. Later, a pulley from the universal base hooks onto the pull loop and pulled in order to achieve flexion of the fingers. The pulley functions also as a differential where it can roll along the bands and distribute the pulling force. In other words, the force is divided to each finger based on the reaction load it experiences. Since the PNG-2 hand is flat, it is 3D printed easily while laying horizontal on the heated bed as demonstrated in Figure 3a.
@@ V Experiments
In the following experiments, we validate the finger model and test the capabilities of various PNG hands with different configurations. As described previously, all hands were 3D printed on a conventional FDM printer with a TPU filament. Videos of the experiments and demonstrations are included in the supplementary material.
@@ V-B Fabrication and Assembly Time
V-C Maximum load
Figures 11 and 12 present the mean maximum loads for grasping rigid objects and for TPU-95A and TPU-85A, respectively. Similarly, Figures 13 and 14 present the mean maximum loads for grasping flexible objects and for TPU-95A and TPU-85A, respectively. The results show that the grippers can withstand significantly high loads roughly proportional to size. The highest load of up to 200 N was achieved by the largest gripper and with rigid objects. When comparing between TPU-95A and TPU-85A, the former can hold higher forces. TPU-95A is much less flexible having a Young modulus of 26MPa compared to TPU-85A with 12MPa. Also, rigid objects are easier to grasp as the flexible object deform between the fingers and slip. While triangular prisms are best grasped, pinch grasps are harder to firmly maintain due to slips.
Another experiment was conducted where the durability of the hand was tested. Different hands were strongly hit with an hammer multiple times followed by grasping experiments. Similarly, the hand was pushed by the robot towards the table and dragged. No damage was caused to the hands and they operated successfully. These experiments can be seen in the supplementary videos.
V-D Grasping Experiments
In the next experiment, we evaluate the ability of PNG hands of various scales printed with TPU-95A to pick-up everyday objects. Note that some of the experiments in this section were not performed with the universal base presented in Section III-D but with earlier prototypes. Nevertheless, the tested PNG hands are in their final form as proposed in this paper. We start by experimenting with five PNG-2 hands having a scale of $\kappa=0.5$, $\kappa=0.75$, $\kappa=1$, $\kappa=1.5$ and $\kappa=1.75$. In addition, we include trials with a PNG-2 hand of scale $\kappa=1.75$ having also fingernails. Each hand was mounted on the universal base while the former is fixed on the Kinova Gen3 robot arm. Various everyday objects were placed on a table for the robot to pick up. The robot was lowered down to the object and stopped at some arbitrary height in the vicinity of the object. After closing the PNG hand, the hand was pulled back up. The success rate for lifting each object over 10 trials is reported in Table I. The results show that smaller hands have more difficulties picking-up objects with large circumference such as the bottle or water glass. With larger hands, higher success rates are reached. Figures 15-17 show snapshots of pick-up trials with a PNG-2 hand of a rubber duck, bunch of keys and wet-wipes, respectively. Figure 18 shows a pick-up of a coin with and without the usage of finger-nails. The addition of fingernails evidently assisted in improving success rate in smaller objects such as the keys and coin. Figures 19 and 20 depict the picking-up of various objects with smaller PNG-2 hands of scale $\kappa=0.75$ and $\kappa=0.5$, respectively.
The same experiments are now presented for PNG-3 and PNG-4 hands of scale $\kappa=0.75$ and $\kappa=1.25$. Table II reports the picking-up success rate for all the hands over all objects and compared to PNG-2. Results show that more fingers slightly improves success rate in general and smaller objects in particular. Nevertheless, a coin is almost impossible to pick-up without fingernails. Figure 21 and 22 show snapshot of picking-up a small bio-hazard bag and various other objects, respectively, with a PNG-4 hand of scale $\kappa=1$. In summary, the results show high success rate for picking up various objects of different sizes with different hands. Moreover, it is possible to use hands of one scale (e.g., $\kappa=1.25$) for successfully picking up a large number of object classes.
@@ V-D Grasping Experiments
TABLE I: Success rates of picking-up and relocating various objects with PNG-2 hands of different scale
@@
Object Scale $\kappa=0.5$ $\kappa=0.75$ $\kappa=1$ $\kappa=1.5$ $\kappa=1.75$ $\kappa=1.75$ (%) (%) (%) (%) (%) w/ nails (%) Hammer 0 0 60 100 100 100 Bottle (0.5L) 0 70 90 100 100 100 Water glass 0 0 0 70 100 100 Chewing gum pack 60 80 100 100 100 100 Cheetos bag 10 50 100 100 100 100 Rubber duck 70 90 100 100 100 100 USB flash drive 100 80 80 90 80 100 Toilet paper roll 100 100 100 100 100 100 Same gripper 100 100 90 80 60 80 Bunch of keys on keychain 100 100 90 80 70 100 Wet wipes from pack 100 100 100 100 100 100 Coin 0 20 20 40 10 100
@@
TABLE II: Success rates of picking-up and relocating various objects with two-, three- and four-finger hands
@@
Object Hand $\kappa=0.75$ $\kappa=1.25$ PNG-2 PNG-3 PNG-4 PNG-2 PNG-3 PNG-4 (%) (%) (%) (%) (%) (%) Hammer 0 0 0 100 100 100 Bottle (0.5L) 70 0 100 100 90 100 Water glass 0 0 50 70 100 100 Chewing gum pack 80 90 100 100 100 100 Cheetos bag 50 50 80 100 100 100 Rubber duck 90 100 100 100 100 100 USB flash drive 80 80 90 90 10 40 Toilet paper roll 100 100 100 100 100 100 Same gripper 100 100 100 80 100 100 Bunch of keys on keychain 100 100 100 80 100 100 Wet wipes from pack 100 100 100 100 100 100 Coin 20 0 0 40 0 0
@@ VI Conclusions
The proposed PNG hand enables a low-cost and disposable solution for chemical and medical applications. Once contaminated, the hand can be disposed and easily replaced by 3D printing a new one. Alternatively, since no electric or metal components are in the hand, it can be cleaned or sterilized for re-use. Future work may further explore other filament materials that are specifically resistant to acidity, radiation or high temperatures. Also, reinforcing the filament with metallic materials may be considered to improve durability but require optimal balance with proper hand flexibility. Due to the robust and strong design, the hand could be further used for precise in-hand manipulations as done in previous work [36, 38, 39, 40]. In such a case, each finger would be independently connected to its own actuator. Future work could also consider an algorithm for automated design of a hand given a specific task or set of objects.

### v2

@@ I Introduction
Figure 2: (a) Front and side view of a PNG finger. Internal view of the proximal link is seen in the circle where the two bands are separated by a bulge. Due to its height, the bulge prevents from the contact pad to press onto the bands. (b) Different scales $\kappa$ of a PNG-2 hand. As convention, a hand with a finger length of 52 mm is defined as $\kappa=1$. Correspondingly, the smallest and largest hands shown and evaluated are with finger lengths of 25.6 mm ($\kappa=0.5$) and 89.7 mm ($\kappa=1.75$), respectively. (c) Examples of three- (PNG-3) and four- (PNG-4) finger hands. Both hands are one-shot 3D printed. (d) Universal base for mounting the printed hand. The 3D printed hand is fixed to the base by inserting the mounting pins into the conical grooves and closing with the tightening cap. A reduction sleeve is used for hands of small scale. (e) Fingernail on the distal link assists in picking-up small objects.
@@ II Related work
One-shot 3D printing is not a novel concept. In particular, monolithic mechanisms have existed for quite some time [21, 22]. However, usages are mostly focused on high-precision positioning [23, 24, 25, 26]. Yet, not much prior work exists on one-shot 3D printing of robotic hands. Of the few, recent work have proposed a one-shot printed granular jamming gripper [27]. Another work focused on one-shot fabrication of a finger while embedding tactile sensors and shape memory alloy actuator [28]. A compliant parallel gripper was proposed and designated for extracting solid objects from powder [29]. Similarly, a bio-inspired parallel gripper was proposed where the joint of each finger receives its compliance from a spiral spring [30]. In [31], a computational approach was presented for designing a one-shot 3D printed two-finger gripper with bi-inspired and sliding joints. RetracTip is a sea-anemone-inspired gripper formed by a bi-stable membrane to achieve delicate grasping [32]. The mechanism is fabricated in a multi-material and high-cost 3D printer (e.g., Stratasys). While innovative, the proposed grippers are not tendon-based and cannot offer high capacity lifting and durability.
A more recent work introduced a one-shot 3D printed two-finger gripper which operates on pneumatic pressure [33]. A work most related to ours proposed a one-shot printed hand with an integrated tendon line [34]. Fabrication of the hand requires a high-cost multi jet fusion printer and, due to the thin tendon, cannot provide high-capacity power grasping or scalability. In addition, it does not enable rapid and easy replacement using a universal base. While not one-shot, a recent and intriguing work has proposed the two-finger underactuated InstaGrasp gripper [35]. The gripper has the structure of the OpenHand [12] Model T42 where each finger is composed of nine 3D printed parts of Polylactic Acid (PLA) and TPU. While almost entirely 3D printed and compliant, the hand requires some assembly. Unlike the proposed PNG hand in this paper, these 3D printed robotic hands do not demonstrate high-load grasping, are not scalable and do not offer easy replacement for disposing if contaminated.
@@ III-B Multi-finger hand
A PNG hand is printed with $n$ fingers in one-shot with a standard FDM printer. For simple notation, we denote an $n$-finger PNG hand as PNG-$n$, e.g., PNG-3 denotes a three-finger PNG hand. Thermoplastic Polyurethane (TPU) filament is a flexible and strong material and, thus, used for the 3D printing. TPU has good corrosion resistance such that chemicals have minor effect [41]. Note that TPU has a heat resistance of up to approximately 80C. Hence, the hand can go through multi-use sterilization by thermal disinfection at 71C or Gamma sterilization [42]. Printing with flexible materials such as TPU requires slower motion of the extruder compared to rigid material. However, early attempts to design a hand with a rigid material such as PLA failed due to rapid plastic deformation in flexion. We make a distinction between a two-finger (PNG-2) and a three or more finger ($n>2$) hands described next.
@@ III-B1 PNG-2, Two-finger hands
A PNG-2 hand is seen in Figure 2b. The tendon bands of both fingers are connected at a pull loop. Later, a pulley from the universal base hooks onto the pull loop and pulled in order to achieve flexion of the fingers. The pulley functions also as a differential where it can roll along the bands and distribute the pulling force. In other words, the force is divided to each finger based on the reaction load it experiences. This is demonstrated in Figure 4. Since the PNG-2 hand is flat, it is 3D printed easily while laying horizontal on the heated bed as demonstrated in Figure 3a.
Figure 4: Demonstration of the differential in the PNG-2. In this case, (left) the right finger is manually constrained with a ball and the left finger compensates for the increase in tendon force $f_{in}$. (right) The ball is not centered in the resulted grasp.
@@ V Experiments
In the following experiments, we validate the finger model and test the capabilities of various PNG hands with different configurations. As described previously, all hands were 3D printed on a conventional FDM printer with a TPU filament. A video of the experiments and demonstrations is available on YouTube [^1] Video link: https://youtu.be/dk5-teuzLGE.
@@ V-B Fabrication and Assembly Time
Figure 16: Success rates of picking-up and relocating various objects with PNG-2 hands of different scale
@@
Object Scale $\kappa=0.5$ $\kappa=0.75$ $\kappa=1$ $\kappa=1.5$ $\kappa=1.75$ $\kappa=1.75$ (%) (%) (%) (%) (%) w/ nails (%) Hammer 0 0 60 100 100 100 Bottle (0.5L) 0 70 90 100 100 100 Water glass 0 0 0 70 100 100 Chewing gum pack 60 80 100 100 100 100 Cheetos bag 10 50 100 100 100 100 Rubber duck 70 90 100 100 100 100 USB flash drive 100 80 80 90 80 100 Toilet paper roll 100 100 100 100 100 100 Same gripper 100 100 90 80 60 80 Bunch of keys 100 100 90 80 70 100 Wet wipes 100 100 100 100 100 100 Coin 0 20 20 40 10 100
@@
Figure 17: Success rates of picking-up and relocating various objects with two-, three- and four-finger hands
@@
Object Hand $\kappa=0.75$ $\kappa=1.25$ PNG-2 PNG-3 PNG-4 PNG-2 PNG-3 PNG-4 (%) (%) (%) (%) (%) (%) Hammer 0 0 0 100 100 100 Bottle (0.5L) 70 0 100 100 90 100 Water glass 0 0 50 70 100 100 Chewing gum pack 80 90 100 100 100 100 Cheetos bag 50 50 80 100 100 100 Rubber duck 90 100 100 100 100 100 USB flash drive 80 80 90 90 10 40 Toilet paper roll 100 100 100 100 100 100 Same gripper 100 100 100 80 100 100 Bunch of keys 100 100 100 80 100 100 Wet wipes 100 100 100 100 100 100 Coin 20 0 0 40 0 0
@@
V-C Maximum load
Figures 12 and 13 present the mean maximum loads for grasping rigid objects and for TPU-95A and TPU-85A, respectively. Similarly, Figures 14 and 15 present the mean maximum loads for grasping flexible objects and for TPU-95A and TPU-85A, respectively. The results show that the grippers can withstand significantly high loads roughly proportional to size. The highest load of up to 200 N was achieved by the largest gripper and with rigid objects. When comparing between TPU-95A and TPU-85A, the former can hold higher forces. TPU-95A is much less flexible having a Young modulus of 26 MPa compared to TPU-85A with 12MPa. Also, rigid objects are easier to grasp as the flexible object deform between the fingers and slip. While triangular prisms are best grasped, pinch grasps are harder to firmly maintain due to slips.
V-D Mechanical characteristics analysis
We next perform a fatigue analysis and test the ability of the PNG hand to withstand cyclic loading. A newly printed PNG-2 hand with $\kappa=1.5$ was set to repeatedly grasp a drill handle of diameter 45 mm. Within each grasp, the hand was closed tight with a tendon force of approximately $f_{in}\approx 75N$. Following every 1,000 trial cycles, the hand was disassembled and the structural integrity was visually inspected. In addition to visual inspection, the hand was tested in loading as performed in Section V-C. The experiment was run for 10,000 cycles. No plastic deformations or any damages were observed but only some wear on the distal contact pads. In addition, loading tests yielded similar results to the ones presented in Figure 12. Overall, the PNG hand is validated to withstand substantial cyclic loading.
In addition to fatigue testing, another experiment was conducted to test the durability of the hand. Different hands were strongly hit with an hammer multiple times followed by grasping experiments. Similarly, the hand was pushed by the robot toward the table and dragged. No damage was caused to the hands and they operated successfully. These experiments can be seen in the supplementary videos.
V-E Grasping Experiments
In the next experiment, we evaluate the ability of PNG hands of various scales printed with TPU-95A to pick-up everyday objects. Note that some of the experiments in this section were not performed with the universal base presented in Section III-D but with earlier prototypes. Nevertheless, the tested PNG hands are in their final form as proposed in this paper. We start by experimenting with five PNG-2 hands having a scale of $\kappa=0.5$, $\kappa=0.75$, $\kappa=1$, $\kappa=1.5$ and $\kappa=1.75$. In addition, we include trials with a PNG-2 hand of scale $\kappa=1.75$ having also fingernails. Each hand was mounted on the universal base while the former is fixed on the Kinova Gen3 robot arm. Various everyday objects were placed on a table for the robot to pick up. The robot was lowered down to the object and stopped at some arbitrary height in the vicinity of the object. After closing the PNG hand, the hand was pulled back up. The success rate for lifting each object over 10 trials is reported in Table 17. The results show that smaller hands have more difficulties picking-up objects with large circumference such as the bottle or water glass. With larger hands, higher success rates are reached. Figures 18-20 show snapshots of pick-up trials with a PNG-2 hand of a rubber duck, bunch of keys and wet-wipes, respectively. Figure 21 shows a pick-up of a coin with and without the usage of finger-nails. The addition of fingernails evidently assisted in improving success rate in smaller objects such as the keys and coin. Figures 23 and 23 depict the picking-up of various objects with smaller PNG-2 hands of scale $\kappa=0.75$ and $\kappa=0.5$, respectively.
The same experiments are now presented for PNG-3 and PNG-4 hands of scale $\kappa=0.75$ and $\kappa=1.25$. Table 17 reports the picking-up success rate for all the hands over all objects and compared to PNG-2. Results show that more fingers slightly improves success rate in general and smaller objects in particular. Nevertheless, a coin is almost impossible to pick-up without fingernails. Figure 24 and 26 show snapshot of picking-up a small bio-hazard bag and various other objects, respectively, with a PNG-4 hand of scale $\kappa=1$. Furthermore, and to demonstrate the ability of the PNG hand to work in complex environments, we also test the grasping of a chewing gum pack in a bucket of liquid with additional objects as seen in Figure 26. The success rate out of ten attempts is 90% with one slippage. In addition, the hand was able to further work in other tasks after getting wet. In summary, the results show high success rate for picking up various objects of different sizes with different hands. Moreover, it is possible to use hands of one scale (e.g., $\kappa=1.25$) for successfully picking up a large number of object classes.
@@ V-E Grasping Experiments
Figure 26: Snapshots of pick-up experiments of a chewing gum pack from a bucket of water with a PNG-2 hand of $\kappa=1.75$.
@@ VI Conclusions
The proposed PNG hand enables a low-cost and disposable solution for chemical and medical applications. Once contaminated, the hand can be disposed and easily replaced by 3D printing a new one. Alternatively, since no electric or metal components are in the hand, it can be cleaned or sterilized for re-use. Future work may further explore other filament materials that are specifically resistant to acidity, radiation or high temperatures. Also, reinforcing the filament with metallic materials may be considered to improve durability but require optimal balance with proper hand flexibility. To speed up printing time, printing with a rigid material can be considered while incorporating flexibility through a unique joint geometry. Due to the robust and strong design, the hand could be further used for precise in-hand manipulations as done in previous work [45, 47, 48, 49]. In such a case, each finger would be independently connected to its own actuator. Future work could also consider an algorithm for automated design of a hand given a specific task or set of objects.